4. Prove that mean of all sample means is an unbiased estimator of population mean by using a random sampling process (n = 2) from a population size of 4 as defined in the following example:
N=4
a=1
b=2
c=3
d=4
Answer #1


Question - population size Population mean = y = 1+2+3+4 = 2.5 Population values are 1,2,3, 4 · N = 4 1+2+3+4 4 All possibleElj) = J, tyrtystya týstys 6 1.5+ 2+2 5+2 5+3+3.5 6 olu = 2.5 6 ie Elj) = Answer Hence Mean of all samble Means is an unbiase